Jennifer is 6 years older than Sue. In 4 years, she will be twice as old as Sue was 5 years ago. Find their ages now.
Answer by LynnMomo(10)

Let Sue's age be x
Then,Jennifer's age be x+6
In 4 years:
(X+6+4)÷2= x-5
(X+10)÷2= x-5
X+10= 2×(x-5)
X+10=2x-10
X-2x=-10-10
-x= -20
X= 20
Sue's age now is 20
Jennifer's age now is 20+6=26.
Gook luck on your study.^_^